# Direct ab initio calculation of magnons in altermagnets: method, spin-space symmetry aspects, and application to MnTe

### v1

@@ IV.1 SSG groups
The standard tool for the analysis of the symmetry properties of the crystalline materials is the apparatus of space groups. For magnetic systems an antisymmetry operation is included into consideration making extension from space groups to magnetic space groupsShubnikov1964 . The antisymmetry operation that in an abstract treatment can be considered as changing the color between black and white in the magnetic case is responsible for the reversal of the magnetization. However, in the problems where the influence of the SOC can be neglected this tool is not sufficient for the description of the properties of magnetic systems. In particular, these limitations have been revealed for both electron band structure calculations and the studies based on the Heisenberg Hamiltonian of interacting atomic moments.
@@
The success of the application of SSG compared with the application of space groups can be characterized as follows. The neglect of the SOC leads effectively to the replacement of the actual physical 3D space by the 6D space where spin and orbital variables are independent and can be transformed separately. The account for this freedom gives important new information about the properties of the system. In addition, the SSGs allow a straightforward establishment of continuity relations between theoretical results obtained within different approximations, such as with and without SOC, or between the results obtained for collinear and helical magnetic configurations. The reason for this is the property that the SSG of the less symmetric case is the subgroup of the SSG of the more symmetric case. The latter feature plays important role in Sec. IV.8 where we discuss the transformation path of the electron band structure of a collinear magnet to the electron band structure of a magnon with a given wave vector q.
@@ IV.2 Degeneracy of the electron states in collinear magnets
On the other hand, in an AFM there must be degeneracy between the states with opposite spin projections. This degeneracy is the reason for the zero net magnetization of the AFM. To expose the origin of the spin degeneracy we first notice that in the Schrödinger equation for electrons with a given spin projection $\sigma$ the electrons see different potentials at the atoms of different magnetic sublattices. Therefore at the first step only the operations leaving the sublattices invariant are considered. For each vector $\bf{k}$ they give star $\{{\bf k}\}_{subl}$ corresponding to the symmetry group of the sublattices. Next it is necessary to specify the symmetry operation that transforms the equations corresponding to different spin projections into each other. The real-space part of this operation $[\alpha^{tr}_{R}|\bm{\tau}]$ must transform the sublattices into each other and be accompanied by an ’antisymmetry’ operation $E^{\prime}$ reversing the signs of the spin indices. As a consequence of such symmetry operation any spin-up state at point $\bf{k}$ is degenerate with a spin-down state at $\alpha^{tr}_{R}\bf{k}$. This makes the system as a whole spin-compensated. The answer to the question whether the spin compensation takes place at each $\bf{k}$ point of the BZ or only between different $\bf{k}$ points depends on the properties of operation $[\alpha^{tr}_{R}|\bm{\tau}]$ transforming the sublattices into each other. If vector $\alpha^{tr}_{R}\bf{k}$ belongs to star $\{{\bf k}\}_{subl}$ the degeneracy takes place at each $\bf{k}$$\in$$\{{\bf k}\}_{subl}$. In the opposite case we deal with an altermagnet with spin splitting at all points of the star $\{{\bf k}\}_{subl}$.
@@ IV.3 Collinear AFM ground state of MnTe
In MnTe, the space group of the atomic lattice contains 24 point operations listed in Table LABEL:table1: 12 operations of type I transform the Mn sublattices into themselves whereas other 12 operations of type II transform them into each other. In the AFM state, the point operations of type II remain the part of the symmetry operations in combination with spin-index reversal operation.
@@ IV.4 Magnetic structure of sublattices in magnon states
Let us discuss the properties of the structure of the AFM magnons that follow from the symmetry arguments. We consider the magnon with an arbitrary wave vector q. The magnetic structure of the magnon is invariant with respect to the generalized translations corresponding to given q. As stated above, the direct magnon calculation includes an effective magnetic field acting on both magnetic sublattices. This field is collinear to the $z$ axis and therefore parallel to the magnetization of one sublattice and antiparallel to the magnetization of the other sublatticecomment_sign_h . Hence the atoms of the two sublattices in the magnon state become inequivalent. On the basis of these conclusions, the magnetic configurations of the two sublattices in a magnon state can be written in the following form
@@
The presence of the ’hidden’ symmetry operation effectively connecting inequivalent Mn sublattices influences also the properties of the Te atoms in the magnon states. In contrast to the Mn sublattices, the atoms of the two Te sublattices remain equivalent because the space inversion transforms the Te sublattices into each other. This operation influences also the directions of the induced Te moments since to keep this symmetry operation intact the induced atomic moments of the two Te sublattices must transform into each other. The straightforward symmetry analysis gives for the Te atoms in the unit cell
@@
where lattice vector ${\bf R}_{IC}$=${\bf a}_{D}$--$I{\bf a}_{C}$. Although the values of angles $\theta_{C},\theta_{D}$ and $\phi_{C},\phi_{D}$ cannot be determined by symmetry arguments the knowledge of relations (13),(14) allows to accelerate the convergence process for the magnon configurations by adequate preparation of the initial magnetic configurations.
@@ IV.7 Instability of the energy of magnon state with respect to the number of magnons
Studying the performance of the method in different situations we applied LDA+$U$ method with various parameters. We noticed that in some cases there is a strong dependence of the magnon-state energy on the small variation of the parameters of the calculations such as the value of the constrained angle $\theta$ or the value of Hubbard parameter $U$. Since the nonmagnetic atoms, in our case Te atoms, play crucial role in the formation of altermagnetism, the character of the hybridization of the electron states of magnetic and nonmagnetic atoms is decisive for the strength of the altermagnetic effects. To influence the position of the Te 5p states with respect to the Mn 3d states we used not only parameter $U_{Mn}$ for the Mn 3d states, usual for the LDA+$U$ method, but also parameter $U_{Te}$ for the Te 5p states. Parameter $U_{Te}$ was treated as a free parameter of the model and was allowed to assume negative values. We emphasize that these model calculations were performed to investigate the potential of the method and we do not connect the results of the modeling with properties of the particular material.
In Fig.12, we show a typical instability behavior. This result was obtained for the energy of the magnon states of type A with wave vector q=(0.1,0.2,0.3) calculated as a function of constrained angle $\theta_{A}$ with parameters $U_{Mn}$=0.2 Ry, and $U_{Te}$=--0.03 Ry. The increas of $\theta_{A}$ can be associated with increasing number of magnons. Around $\theta_{A}$ of 20 the system experiences sharp transformation between two different states. It is straightforward to suppose that in this $\theta_{A}$ region, the increase of $\theta_{A}$ leads to the change of the relative energy positions of different groups of electron states in the band structure of the system. In the iterations, the consequences of this process are enhanced leading to the sharp variation of the energy of the magnon state.
A simular unstable behavior we obtained for magnon B as a function of the constrained angle $\theta_{B}$ (not shown). However, the region of instability was shifted to somewhat smaller angles. As the result, in the instability region we obtained an enhancement of the chirality splitting effect. Thus, for $\theta_{A}$=$\theta_{B}$=20 the difference in energies of the A and B magnon states is $\sim$0.6 mRy that is about two orders of value larger than for angles of 15. Simular instabilty was obtained in calculations for other wave vectors q. Also, this type of singular behavior was obtained for the magnon state energy as the function of parameter $U_{Te}$ (not shown). In the latter calculations constrained angle $\theta$ was fixed at 20. The presence of such instabilities cannot be noticed applying the standard procedure of mapping on the Heisenberg Hamiltonian.
@@
Both chirality splitting and chirality degeneracy of two magnons manifest certain relation between the energies of the magnons. Whether two magnons are equivalent and have equal energies or are inequivalent and have different energies must be reflected in the properties of their electron band structures. In this section we gain an insight into the connection between electronic band structures of the magnon states and chirality properties of these states.
@@ V Conclusions
In the paper we suggest the method for direct ab initio calculation of magnons in complex collinear magnets with weak influence of the spin-orbit coupling. The method does not include the mapping of the electron system on the Heisenberg Hamiltonian of the interacting atomic moments as an intermediate step.
@@
We investigate the connection between the chirality properties of the magnons and the properties of the electron band structures of the magnons. The deep connection between these two different energy characteristics is exposed.
The altermagnetism of MnTe is the consequence of the presence of the Te atoms. Therefore an adequate attention is devoted to the analysis of the selfconsistent response of the Te atoms in the magnon states. In particular, we show that the Te atoms remain equivalent in the magnon states though the Mn atoms beloning to different sublattices became inequivalent. The information about induced magnetic moments of the Te atoms obtained in the symmetry analysis helps to accelerate the convergence of the magnon states by means of optimizing the initial magnetic configuration.
@@
The suggested method accompanied with symmetry arguments on the basis of the spin-space groups provides a new efficient tool for first principles study of the magnons in complex collinear magnets including emerging class of altermagnets.
@@ VI Acknowledgements
K.C. acknowledges financial support by Czech Science Foundation, grant no. 23-­04746S. V.M.S. acknowledges financial support by grant PID2022-139230NB-I00 funded by MCIN/AEI/10.13039/501100011033.

### v2

@@ IV.1 SSG groups
The standard tool for the analysis of the symmetry properties of the crystalline materials is the apparatus of space groups. For magnetic systems an antisymmetry operation is included into consideration making extension from space groups to magnetic space groupsShubnikov1964 ; Padmanabhan2020 . The antisymmetry operation that in an abstract treatment can be considered as changing the color between black and white in the magnetic case is responsible for the reversal of the magnetization direction by means of time reversal. However, in the problems where the influence of the SOC can be neglected this tool is not sufficient for the description of the properties of magnetic systems. In particular, these limitations have been revealed for both electron band structure calculations and the studies based on the Heisenberg Hamiltonian of interacting atomic moments.
@@
The effectiveness of applying SSG, compared to space groups, can be characterized as follows. The neglect of the SOC leads effectively to the replacement of the actual physical 3D space by the 6D space where spin and orbital variables are independent and can be transformed separately. The account for this freedom gives important new information about the properties of the system. In addition, the SSGs allow a straightforward establishment of continuity relations between theoretical results obtained within different approximations, such as with and without SOC, or between the results obtained for collinear and helical magnetic configurations. The reason for this is the property that the SSG of the less symmetric case is the subgroup of the SSG of the more symmetric case. The latter feature plays important role in Sec. IV.8 where we discuss the transformation path of the electron band structure of a collinear magnet to the electron band structure of a magnon with a given wave vector q.
@@ IV.2 Degeneracy of the electron states in collinear magnets
The material of this section contributes to making the paper reasonably self-contained and provides the basis for the comparison with the results on the magnon chirality splitting discussed in Sec. IV.5.
@@
On the other hand, in an AFM there must be degeneracy between the states with opposite spin projections. This degeneracy is the reason for the zero net magnetization of the AFM. To expose the origin of the spin degeneracy we first notice that in the Schrödinger equation for electrons with a given spin projection $\sigma$ the electrons see different potentials at the atoms of different magnetic sublattices. Therefore at the first step only the operations leaving the sublattices invariant are considered. For each vector $\bf{k}$ they give star $\{{\bf k}\}_{subl}$ corresponding to the symmetry group of the sublattices. Next it is necessary to specify the symmetry operation that transforms the equations corresponding to different spin projections into each other. The real-space part of this operation $[\alpha^{tr}_{R}|\bm{\tau}]$ must transform the sublattices into each other and be accompanied by an ’antisymmetry’ operation $E^{\prime}$ reversing the signs of the spin indices. The physical nature of this antisymmetry operation is time reversal. As a consequence of such symmetry operation any spin-up state at point $\bf{k}$ is degenerate with a spin-down state at $\alpha^{tr}_{R}\bf{k}$. This makes the system as a whole spin-compensated. The answer to the question whether the spin compensation takes place at each $\bf{k}$ point of the BZ or only between different $\bf{k}$ points depends on the properties of operation $[\alpha^{tr}_{R}|\bm{\tau}]$ transforming the sublattices into each other. If vector $\alpha^{tr}_{R}\bf{k}$ belongs to star $\{{\bf k}\}_{subl}$ the degeneracy takes place at each $\bf{k}$$\in$$\{{\bf k}\}_{subl}$. In the opposite case we deal with an altermagnet with spin splitting at all points of the star $\{{\bf k}\}_{subl}$.
@@ IV.3 Collinear AFM ground state of MnTe
In MnTe, the space group of the atomic lattice contains 24 point operations listed in Table LABEL:table1: 12 operations of type I transform the Mn sublattices into themselves whereas other 12 operations of type II transform them into each other. In the AFM state, the point operations of type II remain the part of the symmetry operations in combination with spin-index reversal operation (time reversal operation).
@@ IV.4 Magnetic structure of sublattices in magnon states
Let us discuss the properties of the structure of the AFM magnons that follow from the symmetry arguments. These results provide important general knowledge about properties of magnons. Simultaneously, they can be used in setting up and control of the DFT-based calculations. For example, the correlation in the directions of the atomic moments of different atoms predicted by symmetry must be present in the self-consistent solution for the magnon state that provides a tool for controlling the calculations. On the other hand, such a correlation can be used from the beginning of calculations to decrease the number of degrees of freedom and simplify the process of convergence. In addition, the symmetry-based knowledge that the magnetic atoms become inequivalent in the magnon excitations helps to avoid an unphysical assumption of the preserved equivalence of atoms and an erroneous constraint of the equivalence of magnetic atoms in the calculations.
We consider the magnon with an arbitrary wave vector q. The magnetic structure of the magnon is invariant with respect to the generalized translations corresponding to given q. As stated above, the direct magnon calculation includes an effective magnetic field acting on both magnetic sublattices. This field is collinear to the $z$ axis and therefore parallel to the magnetization of one sublattice and antiparallel to the magnetization of the other sublatticecomment_sign_h . Hence the atoms of the two sublattices in the magnon state become inequivalent. On the basis of these conclusions, the magnetic configurations of the two sublattices in a magnon state can be written in the following form
@@
The presence of the symmetry operation effectively connecting inequivalent Mn sublattices influences also the properties of the Te atoms in the magnon states. In contrast to the Mn sublattices, the atoms of the two Te sublattices remain equivalent because the space inversion transforms the Te sublattices into each other. This operation influences also the directions of the induced Te moments since to keep this symmetry operation intact the induced atomic moments of the two Te sublattices must transform into each other. The straightforward symmetry analysis gives for the Te atoms in the unit cell
@@
where lattice vector ${\bf R}_{IC}$=${\bf a}_{D}$--$I{\bf a}_{C}$. Although the values of angles $\theta_{C},\theta_{D}$ and $\phi_{C},\phi_{D}$ cannot be determined by symmetry arguments the knowledge of relations (13),(14) allows to accelerate the convergence process for the magnon configurations by adequate preparation of the initial magnetic configurations. The calculation process includes the self-consistent determination of the directions of all atomic moments. This process can be rather time consuming. The knowledge of the symmetry relations between the moments of different atoms allows to decrease the number of the degrees of freedom and leads to a considerable reduction of the convergence time.
@@ IV.7 Instability of the energy of magnon state with respect to the number of magnons
While studying the performance of the method we carried out numerous model calculations. In particular, we investigated the consequences of the relative energy shift of the Mn 3d and Te 5p states that influences the hybridization between these states. Since the nonmagnetic atoms, in our case Te atoms, play crucial role in the formation of altermagnetism, the character of the hybridization of the electron states of magnetic and nonmagnetic atoms is an important factor influencing the character of the altermagnetic effects. As a convenient tool for varying the energy positions of the Mn 3d and Te 5p states we employed LDA+$U$ method with various values of the corresponding $U$ parameters. In these numerical experiments we used not only parameter $U_{Mn}$ for the Mn 3d states but also small $U_{Te}$ parameter for the Te 5p states, the latter treated as a free parameter assuming both positive and negative values. In these model calculations, we noticed that in some cases the spin wave states behave singular as the function of $\Delta m_{z}$.
In Fig.12, we show a typical instability behavior. This result was obtained for the energy of the magnon states of type A with wave vector q=(0.1,0.2,0.3) calculated as a function of constrained angle $\theta_{A}$ with parameters $U_{Mn}$=0.2 Ry, and $U_{Te}$=--0.03 Ry. The increase of $\theta_{A}$ can be associated with increasing number of magnons. Around $\theta_{A}$ of 20 the system experiences sharp transformation between two different states. It is straightforward to suppose that in this $\theta_{A}$ region, the increase of $\theta_{A}$ leads to the change of the relative energy positions of different groups of electron states in the band structure of the system. In the iterations, the consequences of this process are enhanced leading to the sharp variation of the energy of the magnon state.
A simular unstable behavior we obtained for magnon B as a function of the constrained angle $\theta_{B}$ (not shown). However, the region of instability was shifted to somewhat smaller angles. As the result, in the instability region we obtained an enhancement of the chirality splitting effect. Thus, for $\theta_{A}$=$\theta_{B}$=20 the difference in energies of the A and B magnon states is $\sim$0.6 mRy that is about two orders of value larger than for angles of 15. Simular instabilty was obtained in calculations for other wave vectors q. The presence of such instabilities cannot be noticed applying the standard procedure of mapping on the Heisenberg Hamiltonian. The possibility to study such instabilities that are different for different magnon chirality channels is a useful feature of the suggested method.
@@
Both chirality splitting and chirality degeneracy of two magnons manifest certain relation between the energies of the magnons. Whether two magnons are equivalent and have equal energies or are inequivalent and have different energies must be reflected in the properties of their electron band structures. In this section we gain an insight into the connection between electronic band structures of the magnon states and chirality properties of these states. Our method includes the calculation of the electronic band structures of the magnons as a part of the itiration process.
@@ V Conclusions
In the paper we suggest the method for direct ab initio calculation of magnons in complex collinear magnets without the spin-orbit coupling. The method does not include the mapping of the electron system on the Heisenberg Hamiltonian of the interacting atomic moments as an intermediate step.
@@
Our method allows investigation of the electron band structure of individual magnons. We investigate the connection between the chirality properties of the magnons and the properties of the electron band structures of the magnons. The deep connection between these two different energy characteristics is exposed.
The altermagnetism of MnTe is the consequence of the presence of the Te atoms. The method allows the study of the self-consistent response of the nonmagnetic atoms to the formation of magnon excitations. An adequate attention is devoted to the analysis of the properties of the Te atoms in magnons in MnTe. In particular, we show that the Te atoms remain equivalent in the magnon states though the Mn atoms beloning to different sublattices became inequivalent. The information about induced magnetic moments of the Te atoms obtained in the symmetry analysis besides providing important physical information helps to accelerate the numerical convergence of the magnon states by means of decreasing the number of degrees of freedom in the self-consistent calculations.
@@
The suggested method accompanied with symmetry arguments on the basis of the spin-space groups provides a new tool making possible efficient direct first principles study of the magnons in complex collinear magnets including emerging class of altermagnets.
@@
An important question is if the method can be applied to the systems where the SOC cannot be neglected. The main problem with the SOC is that it reduces the 6D space with separate spin and orbital variables to the actual 3D space where any point transformation acts on both types of variables. In the presence of SOC, the generalized translations do not commute with the Kohn-Sham Hamiltonian and the reduction of the problem, for an arbitrary wave vector q, to the small crystallographic unit cell cannot be performed. This mathematical problem is the reflection of the physical reality consisting in the fact that the SOC disturbs the perfect helical structure of the spin waves by means of the magnetic anisotropy. However, there is an important class of materials where the method can potentially be applied in the presence of the SOC. These are the materials with uniaxial magnetic anisotropy. If the $z$ axis is an easy axis and the anisotropy in the $xy$-plane is negligible, the spin rotation by an arbitrary angle about the anisotropy axis leaves the system invariant, and the generalized translation symmetry remains intact. In the Ref. Sandratskii2017, , the possibility to neglect the in-plane anisotropy was used to study the Dzyaloshinskii-Moriya interaction that is one of the consequences of SOC. The extension of the direct method suggested in this paper to the case of the magnets with uniaxial anisotropy is an interesting problem for the future studies. Another possibility to deal with the SOC is, first, to perform self-consistent calculation of the magnon states without SOC and then to take into account the SOC within the first-order perturbation theory Heide2009 .
@@ VI Acknowledgements
The discussions with Samir Lounis are gratefully acknowledged. K.C. acknowledges financial support by Czech Science Foundation, grant no. 23-04746S, and project Quantum materials for applications in sustainable technologies (QM4ST), funded as project no. CZ.02.01.01/00/22_008/0004572 by P JAK, Ministry of Education, Youth and Sports of the Czech Republic. V.M.S. acknowledges financial support by grant PID2022-139230NB-I00 funded by MCIN/AEI/10.13039/501100011033.